Find an equation of the line passing (2, 1) having a slope of -3
Answer by ankor@dixie-net.com(22740)

Find an equation of the line passing (2, 1) having a slope of -3
:
Use the point slope form: y - y1 = m(x - x1)
x1 = 2; y1 = 1; m =-3
:
y - 1 = -3(x - 2)
y - 1 = -3x + 6
y = -3x + 6 + 1
y = -3x + 7
:
:
You can confirm this, replace x with 2, ensure it equals 1